OpenStudy (anonymous):

2. Let f be the function shown below, with domain the closed interval [0,6]. The graph of f is shown below, composed of two semicircles.

OpenStudy (anonymous):

@agent0smith it's suppose to say "find the domain of..."

OpenStudy (anonymous):

i want to know if we use the equation of a circle and divide it by 2, or the area of a circle and divide it by 2...

OpenStudy (agent0smith):

domain of f(t) is only from 0 to 6, so the domain of h(x) is the same, except it's from 0 to 2x-1, where 2x-1 must equal 6. Find x.

OpenStudy (anonymous):

is t the independent variable on the graph? it's kinda fuzzy

OpenStudy (agent0smith):

Yeah idk why the question looks so terrible, crap all over the place.

OpenStudy (anonymous):

well, then it's 0 <= 2x - 1 <= 6

OpenStudy (anonymous):

so do main of h(x) is 1/2 <= x < 7/2

OpenStudy (anonymous):

where does the integral come in place?

OpenStudy (anonymous):

0 to x is 3

OpenStudy (anonymous):

well, h(x) is defined by taking the integral of f(t), for t in [0,6]. There and since 2x-1 is in that interval, 0 <= 2x-1 <= 6

OpenStudy (anonymous):

i mean x=3, so the domain is [0,3]

OpenStudy (anonymous):

sorry about the pic, you guys.

OpenStudy (anonymous):

when when x = 3,